# Bifurcations of the Hénon map with additive bounded noise

### v1

@@ 4.3. Topological bifurcation due to a heteroclinic bifurcation of the boundary map
Consider the Hénon map with bounded noise (1.1) at parameter values $b=0.3$ and $\varepsilon=0.6$ and $a=0.36$. This random dynamical system has a unique minimal attractor, and it is displayed together with its domain of attraction in Figure 12(a). When $a$ is increased, the attractor moves towards its dual repeller (cf. Figure 12(b)), and after it collides, when $a\approx 0.37427117$, the attractor disappears instantaneously. From the set-valued point of view, this scenario is similar to that of the example in Section 4.2, above, cf. also [lamb2015topological].
@@
Figure 12. Relevant projections of parts of invariant manifolds involved in the topological bifurcation of the minimal attractor of the Hénon map with bounded noise (1.1) at parameter values $a\approx 0.37427117$, $b=0.3$, and $\varepsilon=0.6$: the unstable manifold (blue curve) of a saddle fixed point (red dot) for the boundary of the attractor and the stable manifold (orange curve) of another saddle fixed point (purple dot) for the boundary of its dual repeller. In (a), at $a=0.37$, we observe a situation close to a heteroclinic connection between the invariant manifolds (the intersection of these manifolds at $a\approx 0.37427117$ in the three-dimensional state space of the boundary map manifests itself as a tangency between the corresponding projections), compare with Figure 12(b). At $a=0.38$, after the topological bifurcation has taken place and the attractor has exploded, the projection of the unstable manifold (blue) of the saddle fixed point (red) appears to be in (b).
@@
Figure 13. The boundary of the minimal attractor and its dual repeller represented by (parts of) relevant projections of invariant manifolds at the parameter values: $b=0.3,\varepsilon=0.6,a=0.37427117$, on the verge of a topological bifurcation. The projections to the interior of the minimal attractor or its dual repeller are omitted for clarity. A cascade of wedge singularities is found on both sides of the boundary of the dual repeller near the saddle fixed point (red dot in (a)). There is also a cascade of wedge singularities on both sides of the boundary of the minimal attractor near the other saddle fixed point (purple dot in (a) and the magnification (b), where tangencies are visible only on one side of the saddle point due to the small scale at which they arise on the other side). In [lamb2020boundaries], boundary points with cascades of wedges on both sides are called shallow-shallow singularities.
In Figure 12(a), the projection of the unstable manifold (in blue) of the red fixed point appears to be on the verge of intersecting with the projection of the stable manifold (in orange) of the purple fixed point. Indeed, when $a$ is increased further, at $a\approx 0.37427117$ the unstable manifold forms a heteroclinic cycle with the stable manifold (also cf. Figure 13). The birth of this heteroclinic cycle of the boundary map corresponds to a topological bifurcation of the minimal attractor, where the minimal attractor disappears after the bifurcation. The heteroclinic intersections represent collision points between the minimal invariant set and its dual repeller (when the minimal invariant set loses its attractivity). This is a known necessary condition for topological bifurcation [lamb2015topological].
We present a continuation of the saddle fixed points and their respective one-dimensional unstable and stable manifolds for $a=0.38$ in Figure 12(b): the heteroclinic cycle between the unstable manifold of the red fixed point and the stable manifold of the purple fixed point is broken, leading to the topological bifurcation. The unstable manifold of the red fixed point appears to be (towards the left edge of Figure 12(b)). This signifies that the minimal attractor from $a=0.37$ (in Figure 12(a)) explodes, and has disappeared after the bifurcation.
@@
We observe a two-sided cascade of wedge singularities on the boundary of the dual repeller (in orange), near the projection of the saddle fixed point (in red). There is also a two-sided cascade of wedge singularities on the boundary of the minimal attractor (in blue), near the projection of the saddle fixed point (in purple). On the magnification in Figure 13(b), the heteroclinic tangencies are only visible on one side of the saddle point as they arise on a very small scale on the other side. This points to the existence of a shallow-shallow singularity [lamb2020boundaries] on both the boundary of the dual repeller and the minimal invariant set, observed only at the point of topological bifurcation (where the invariant set loses its attractivity). The genericity of this kind of topological bifurcation, in relation to the underlying geometry, requires further theoretical explanation, which is beyond the scope of this paper.
@@
The existence and persistence of such a cascade may be understood from the point of view of the boundary map. In Figure 15, we present a visualisation of the unstable manifold (blue) of a fixed point of the boundary map in $\mathbb{R}^{2}\times S^{1}$, the projection of which (red) yields Figure 14(d). As the unstable manifold approaches the stable periodic point, it ’wiggles’. This wiggling appears to persist arbitrarily close to the periodic point. Such infinite wiggling indeed arises persistently if the unstable manifold of the fixed point has a non-transversal intersection with the strong stable foliation of the stable two-periodic orbit, as illustrated schematically in Figure 16. It appears that due to the underlying geometry of the boundary map, the relevant projection of each such a wiggle leads to a self-intersection and thus to a wedge singularity on the boundary, leading to an infinite cascade of wedge singularities accumulating to the projection of the periodic point. A formal proof requires a deep appreciation of the underlying geometry, which is beyond the scope of this paper. The possibility of such a cascade was already noted in [lamb2020boundaries], where the accumulation point was coined a shallow singularity. Interestingly, the dynamical setting indicates that such accumulation points of singularities can arise persistently, whereas this would not be expected from singularity theoretical considerations of projections of smooth curves.
@@
However, we observe that the disappearance of the shallow singularity singularities is characterised by a scaling relation when the parameter $a$ is gradually decreased towards the bifurcation point from 0.49 to around 0.44. We observe the cascade disappears by a mechanics of successive boundary bifurcations where, one by one, the outermost wedge singularity of the cascade, furthest from the shallow singularity, disappears.
@@ 5. Outlook
In this paper, we have employed the boundary map (3.1) (first introduced in [kourliouros2023persistence]) as a tool to numerically study attractors of random dynamical systems with bounded noise and their bifurcations. The Hénon map with bounded noise (1.1) was chosen as a prototypical example, for illustrative purposes.
@@
As a first step, [kourliouros2023persistence] addresses the smooth persistence of minimal invariant sets with smooth boundaries, using normal hyperbolicity and relevant insights from contact geometry (that underlies the setting). Extensions of these results are expected to yield a classification of attractors that are persistent in terms of topology and boundary singularity structure. Indeed, the present paper contains various conjectures in this direction, supported by numerics and intuitive arguments.
@@ Acknowledgments
$h_{f}(x,n)=\left(f(x),\frac{(f^{\prime}(x)^{T})^{-1}n}{||(f^{\prime}(x)^{T})^{-1}n||}\right),\quad\mbox{and}\quad g_{\varepsilon}(x,n)=(x+\varepsilon n,n).$
@@ Appendix B Properties of eigenvalues and eigenvectors of the derivative at fixed points of the boundary map
We note the properties highlighted in Proposition B.1 also hold for periodic points, see [tey2022minimal, Chapter 5].

### v2

@@ 4.3. Topological bifurcation due to a heteroclinic bifurcation of the boundary map
Consider the Hénon map with bounded noise (1.1) at parameter values $b=0.3$ and $\varepsilon=0.6$ and $a=0.36$. This random dynamical system has a unique minimal attractor, and it is displayed together with its domain of attraction in Figure 12(a). When $a$ is increased, the attractor moves towards its dual repeller (cf. Figure 12(b)), and after it collides, when $a\approx 0.37427117$, the attractor disappears instantaneously. From the set-valued point of view, this scenario is similar to that of the example in Section 4.2, above, cf. also [12].
@@
Figure 12. Relevant projections of parts of invariant manifolds involved in the topological bifurcation of the minimal attractor of the Hénon map with bounded noise (1.1) at parameter values $a\approx 0.37427117$, $b=0.3$, and $\varepsilon=0.6$: the unstable manifold (blue curve) of a saddle fixed point (red dot) for the boundary of the attractor and the stable manifold (orange curve) of another saddle fixed point (purple dot) for the boundary of its dual repeller. In (a), at $a=0.37$, we observe a situation close to a heteroclinic connection between the invariant manifolds (the intersection of these manifolds at $a\approx 0.37427117$ in the three-dimensional state space of the boundary map manifests itself as a tangency between the corresponding projections), compare with Figure 12(b). At $a=0.38$, after the topological bifurcation has taken place and the attractor has exploded, the projection of the unstable manifold (blue) of the saddle fixed point (red) appears to be unbounded in (b).
@@
Figure 13. The boundary of the minimal attractor and its dual repeller represented by (parts of) relevant projections of invariant manifolds at the parameter values: $b=0.3,\varepsilon=0.6,a=0.37427117$, on the verge of a topological bifurcation. The projections to the interior of the minimal attractor or its dual repeller are omitted for clarity. A cascade of wedge singularities is found on both sides of the boundary of the dual repeller near the saddle fixed point (red dot in (a)). There is also a cascade of wedge singularities on both sides of the boundary of the minimal attractor near the other saddle fixed point (purple dot in (a) and the magnification (b), where tangencies are visible only on one side of the saddle point due to the small scale at which they arise on the other side). In [13], boundary points with cascades of wedges on both sides are called shallow-shallow singularities.
In Figure 12(a), the projection of the unstable manifold (in blue) of the red fixed point appears to be on the verge of intersecting with the projection of the stable manifold (in orange) of the purple fixed point. Indeed, when $a$ is increased further, at $a\approx 0.37427117$ the unstable manifold forms a heteroclinic cycle with the stable manifold (also cf. Figure 13). The birth of this heteroclinic cycle of the boundary map corresponds to a topological bifurcation of the minimal attractor, where the minimal attractor disappears after the bifurcation. The heteroclinic intersections represent collision points between the minimal invariant set and its dual repeller (when the minimal invariant set loses its attractivity). This is a known necessary condition for topological bifurcation [12].
We present a continuation of the saddle fixed points and their respective one-dimensional unstable and stable manifolds for $a=0.38$ in Figure 12(b): the heteroclinic cycle between the unstable manifold of the red fixed point and the stable manifold of the purple fixed point is broken, leading to the topological bifurcation. The unstable manifold of the red fixed point appears to be unbounded (towards the left edge of Figure 12(b)). This signifies that the minimal attractor from $a=0.37$ (in Figure 12(a)) explodes, and has disappeared after the bifurcation.
@@
We observe a two-sided cascade of wedge singularities on the boundary of the dual repeller (in orange), near the projection of the saddle fixed point (in red). There is also a two-sided cascade of wedge singularities on the boundary of the minimal attractor (in blue), near the projection of the saddle fixed point (in purple). On the magnification in Figure 13(b), the heteroclinic tangencies are only visible on one side of the saddle point as they arise on a very small scale on the other side. This points to the existence of a shallow-shallow singularity [13] on both the boundary of the dual repeller and the minimal invariant set, observed only at the point of topological bifurcation (where the invariant set loses its attractivity). The genericity of this kind of topological bifurcation, in relation to the underlying geometry, requires further theoretical explanation, which is beyond the scope of this paper.
@@
The existence and persistence of such a cascade may be understood from the point of view of the boundary map. In Figure 15, we present a visualisation of the unstable manifold (blue) of a fixed point of the boundary map in $\mathbb{R}^{2}\times S^{1}$, the projection of which (red) yields Figure 14(d). As the unstable manifold approaches the stable periodic point, it ‘wiggles’. This wiggling appears to persist arbitrarily close to the periodic point. Such infinite wiggling indeed arises persistently if the unstable manifold of the fixed point has a non-transversal intersection with the strong stable foliation of the stable two-periodic orbit, as illustrated schematically in Figure 16. It appears that due to the underlying geometry of the boundary map, the relevant projection of each such a wiggle leads to a self-intersection and thus to a wedge singularity on the boundary, leading to an infinite cascade of wedge singularities accumulating to the projection of the periodic point. A formal proof requires a deep appreciation of the underlying geometry, which is beyond the scope of this paper. The possibility of such a cascade was already noted in [13], where the accumulation point was coined a shallow singularity. Interestingly, the dynamical setting indicates that such accumulation points of singularities can arise persistently, whereas this would not be expected from singularity theoretical considerations of projections of smooth curves.
@@
However, we observe that the disappearance of the shallow singularity is characterised by a scaling relation when the parameter $a$ is gradually decreased towards the bifurcation point from 0.49 to around 0.44. We observe the cascade disappears by the mechanics of successive boundary bifurcations where, one by one, the outermost wedge singularity of the cascade, furthest from the shallow singularity, disappears.
@@ 5. Outlook
In this paper, we have employed the boundary map (3.1) (first introduced in [11]) as a tool to numerically study attractors of random dynamical systems with bounded noise and their bifurcations. The Hénon map with bounded noise (1.1) was chosen as a prototypical example, for illustrative purposes.
@@
As a first step, [11] addresses the smooth persistence of minimal invariant sets with smooth boundaries, using normal hyperbolicity and relevant insights from contact geometry (that underlies the setting). Extensions of these results are expected to yield a classification of attractors that are persistent in terms of topology and boundary singularity structure. Indeed, the present paper contains various conjectures in this direction, supported by numerics and intuitive arguments.
@@ Acknowledgments
[…]
@@ Appendix B Properties of eigenvalues and eigenvectors of the derivative at fixed points of the boundary map
We note the properties highlighted in Proposition B.1 also hold for periodic points, see [18, Chapter 5].